between two stations a train accelerates uniformly at first then moves with constant speed and finally retards uniformly.if the ratio of time taken are 1:8:1 and the greatest speed is 60km/hr , what is the avg speed over the whole journey ? In this question, why do we take total time as 10t? can anyone help?


Grade:11

1 Answers

Sourav
6 years ago
See the average velocity is the ratio of total distance and total time. Since the ratio of times are 1:8:1. If the proportionality constant is k then t1= k t2= 8k and t3= k so the total time is 10k and the total distance is the area of trapezium = 1/2 (8k+10k)* 60= 540k so the average velocity is 540k/10k = 54 km/hr